A person, living in Canada in the winter, pumps their car tires in the morning with a pressure of 35 psi. The temperature in the morning is 20°C. By the evening, the temperature has dropped to 0°C and the tire pressure monitoring system in the car starts indicating low pressure with a warning light. What pressure will the tires have in the evening that made the warning light from the pressure monitoring system turn on?
1
Expert's answer
2020-10-16T08:57:23-0400

The pressure of a gas "p" is related to its temperature "T", volume "V" and its quantity "n" through the ideal gas law:

"pV=nRT" .

For the gas in the car tires, the volume and the number of the moles of the gas are constant:

"V,n = const" .

Therefore, the ratio of pressure and temperature should be kept constant, too:

"\frac{p}{T} = const".

Remember that the temperature should be experssed in kelvin and not in celsius. Therefore, 20°C is 293.15 K and 0°C is 273.15 K:

"\frac{p_{morning}}{T_{morning}}=\frac{p_{evening}}{T_{evening}}"

"\frac{35}{293.15} = \frac{p_{evening}}{273.15}"

"p_{evening}=\frac{35}{293.15}\cdot273.15=32.6" psi.

Answer: the tires have a pressure of gas of 32.6 psi in the evening.